OpenStudy (anonymous):

Solve for x: -2(5 - 3x) = 12 - 3(5 - x) *7/6 *-7/6 *7/3 *-7/3

OpenStudy (anonymous):

expand the terms

OpenStudy (anonymous):

-2(5-3x)=12-3(5-x) -10+6x=12-15+3x can you do the remaining part?

OpenStudy (anonymous):

do somthing that will make them equal

OpenStudy (anonymous):

I think so! I will give the answer in a sec

OpenStudy (anonymous):

ugh... Im going to guess...

OpenStudy (anonymous):

1 minute has gone already lol :P move those 'x's on the same side, and move the integers on the other side

OpenStudy (anonymous):

-10+6x=12-15+3x 6x-3x=12-15+10 3x=7 then x=?

OpenStudy (anonymous):

7/3

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

yes